OpenStudy (anonymous):

15x^5-20x^4=6x^3-8x^2

Directrix (directrix):

15x^5-20x^4=6x^3-8x^2 15x^5-20x^4-6x^3+8x^2=0 x² ( 15 x³ - 20 x² - 6x + 8) = 0 x² = 0 OR ( 15 x³ - 20 x² - 6x + 8) = 0 (Zero Product Property) x = 0 or 0 (Zero is a double root) OR 15 x³ - 20 x² - 6x + 8 = 0 @wixitodeleon Your Task: Solve 15 x³ - 20 x² - 6x + 8 = 0 for x. Suggestion: Use the Rational Root Theorem to find a root.